# Intent Prediction-Driven Model Predictive Control for UAV Planning and Navigation in Dynamic Environments

### v1

@@ Abstract
The emergence of indoor aerial robots holds significant potential for enhancing construction site workers’ productivity by autonomously performing inspection and mapping tasks. The key challenge to this application is ensuring navigation safety with human workers. While navigation in static environments has been extensively studied, navigating dynamic environments remains open due to challenges in perception and planning. Payload limitations of unmanned aerial vehicles limit them to using cameras with limited fields of view, resulting in unreliable perception and tracking during collision avoidance. Moreover, the unpredictable nature of the dynamic environments can quickly make the generated optimal trajectory outdated. To address these challenges, this paper presents a comprehensive navigation framework that incorporates both perception and planning, introducing the concept of dynamic obstacle intent prediction. Our perception module detects and tracks dynamic obstacles efficiently and handles tracking loss and occlusion during collision avoidance. The proposed intent prediction module employs a Markov Decision Process (MDP) to forecast potential actions of dynamic obstacles with the possible future trajectories. Finally, a novel intent-based planning algorithm, leveraging model predictive control (MPC), is applied to generate safe navigation trajectories. Simulation and physical experiments [^3] Experiment video link: https://youtu.be/UeBShELDzyM demonstrate that our method enables safe navigation in dynamic environments and achieves the fewest collisions compared to benchmarks.
@@ I Introduction
In recent years, lightweight autonomous unmanned aerial vehicles (UAVs) have shown great potential for inspection and mapping in indoor construction sites [1][2][3]. Unlike the open outdoor environments, indoor construction sites present more complex conditions, including the presence of moving workers, machines, and other static structures in close proximity. Ensuring the operational safety of these autonomous robots in relation to human workers is thus a critical challenge that must be addressed before their deployment. As a result, designing a navigation framework that enables safe operation in dynamic environments is essential.
Despite the success of previous research [4][5] in navigation for complex static environments, achieving safe robot navigation in dynamic environments remains challenging due to two major factors. First, the low image quality of onboard cameras in lightweight UAVs makes accurate detection and tracking of dynamic obstacles difficult. Although our previous work [6] utilizes an ensemble detection method to improve accuracy and computational efficiency, the camera’s limited field of view (FOV) can still result in tracking failures during collision avoidance maneuvers, increasing the risk of severe collisions from undetected dynamic obstacles. Second, the unpredictable motion of dynamic obstacles can quickly make an optimized trajectory ineffective. While many previous works [7][8][9][10] have incorporated obstacle trajectory prediction into planning frameworks, most methods consider only a single future mode (i.e., one trajectory) per obstacle, typically using a simple linear motion model. Although these approaches can improve dynamic obstacle avoidance, their effectiveness is significantly diminished if the actual motion of the obstacles differs from the assumed model.
@@
To address these challenges, this paper presents an intent prediction-driven and comprehensive navigation framework based on a model predictive control scheme. An enhanced perception module for dynamic obstacle detection and tracking is introduced to mitigate the lost-tracking issue during collision avoidance maneuvers. Crucially, to enable the planner to consider all potential future movements of dynamic obstacles, this work introduces the concept of dynamic obstacle intents, represented as a probability distribution that reflects the likelihood of predefined high-level decisions for each obstacle. Instead of predicting a single trajectory per obstacle, the prediction module generates all possible trajectories based on various possible intents. The proposed trajectory planning algorithm optimizes multiple trajectories according to potential future states using model predictive control formulation and applies a scoring function to select the best trajectory for execution. Fig. 1 shows an example of our UAV navigating in a dynamic environment with the proposed method. The contributions of this work are:
  • Lost-Tracking-Aware Perception: Building on previous work in obstacle detection and tracking [6], the perception module is enhanced with an out-of-view compensation method to mitigate lost tracking and occlusion during collision avoidance maneuvers.
  • Intent and Trajectory Prediction: This work introduces dynamic obstacle intent prediction, which forecasts potential actions and future trajectories using a Markov Decision Process, enabling more comprehensive anticipation of movements within the environment.
  • Intent-Based Planning Algorithm: Unlike traditional methods optimizing a single trajectory based on one future prediction per obstacle, our algorithm generates multiple candidates and selects the highest-score one by evaluating potential future states, thereby reducing trajectory flickering in unpredictable environments.
@@ II Related Work
Recent years have seen significant advances in lightweight indoor UAV navigation. While early challenges in trajectory planning [11][12] and collision avoidance [13][14] for static environments have been largely addressed [4][5][15], navigating dynamic environments remains a challenge due to the unpredictable nature of dynamic obstacles. Research efforts to improve UAV navigation safety and efficiency in dynamic environments can be broadly categorized into reactive planning-based and predictive planning-based methods.
@@
Reactive planning-based methods: These methods treat dynamic obstacles as static at each time step, adapting static planning methods to dynamic environments through high-frequency replanning. Oleynikova et al. [16] use depth vision to detect obstacles and generate waypoints for collision avoidance. In [4][5], occupancy-based maps are adopted for environment representation, which enables safe trajectory generation. Similarly, Ren et al. [15] achieve aggressive and safe maneuvering using whole-body motion planning. However, because these methods are primarily designed for static obstacles, their lack of dynamic obstacle perception and planning can lead to potential failures. To address the future uncertainties of dynamic obstacles, Guo et al. [17] construct risk regions, while [18] adopts a safe flight corridor that considers both static and dynamic obstacles. In [19], moving obstacle avoidance is achieved through raycasting and Riemannian motion policy. Although some reactive methods can achieve safe navigation, they can become significantly less effective and risky in highly dynamic environments.
@@
Predictive planning-based methods: Different from the reactive methods, methods from this category make plans utilizing the prediction of dynamic obstacles in a future horizon. In [20][7][21], a chance-constrained model predictive control approach is used to account for the uncertainties of dynamic obstacles. Similarly, Jian et al. [22] employ a control barrier function within an MPC framework for collision avoidance. Wang et al. [8] enhance safe navigation by combining occlusion-aware obstacle tracking with spline-based trajectory optimization. Chen et al. [23] sample and evaluate trajectories based on predicted dynamic obstacles, while [10] incorporates UAV yaw angle for collision avoidance. In [9], a vision-aided planner is used for avoiding both static and dynamic obstacles. For dynamic obstacle perception, a learning-based method [24] is proposed for 3D detection and tracking, with further improvements in computation and accuracy achieved through ensemble detection in [6]. Additionally, some methods [25][26] use learning-based approaches to avoid dynamic obstacles. However, these methods only rely on a constant velocity model for predicting the future state of dynamic obstacles, which may not always be reliable. To better model pedestrian trajectories, Peddi et al. [27] trained a Hidden Markov Model on trajectory datasets, and Thomas et al. [28] applied self-supervised learning for map-level dynamic obstacle prediction. Although these methods, along with other popular learning-based approaches [29][30], improve upon traditional linear predictions, they primarily focus on low-level trajectory prediction and fail to account for the high-level decision-making processes of dynamic obstacles, potentially overlooking variations in obstacle behavior. Inspired by methods in autonomous driving [31][32][33][34] that predict multiple future trajectories for vehicles, we introduce the concept of dynamic obstacle intent. Our approach considers the high-level decisions of obstacles, predicts trajectories based on these intents, and generates safe trajectories accordingly.
III Methodology
@@
The proposed method contains three main modules: perception, prediction, and planning, as shown in Fig. 2. The perception module (Sec. III-A) is divided into two parts, handling static and dynamic obstacle perception, respectively. The intent and trajectory prediction module (Sec. III-B) processes dynamic obstacle history from the dynamic perception module and the occupancy map from the static perception module to generate intent probability distributions and corresponding predicted trajectories for dynamic obstacles. The intent-based planning algorithm (Sec. III-D) includes a trajectory planning that uses a model predictive control formulation (Sec. III-C) and a score-based trajectory selection module to produce the final trajectory for execution.
III-A Lost-Tracking-Aware Perception
@@
The perception module consists of a static and a dynamic component for separate obstacle detection. As shown in Fig. 2, the static perception uses a traditional occupancy voxel map to represent static obstacles, with dynamic cleaning to remove noise caused by moving obstacles. The dynamic perception is based on [6] and incorporates an updated ensemble detection method along with out-of-view compensation to improve accuracy and safety during collision avoidance, particularly in dynamic-obstacle-tracking loss scenarios.
@@
Figure 3: Illustration of unsafe dynamic obstacle avoidance due to tracking loss. (a) The robot initially detects and tracks the dynamic obstacle, generating a safe trajectory. (b) As the robot maneuvers, the obstacle moves out of the camera view, leading to an unsafe replanned trajectory.
@@
Dynamic Obstacle Detection: The main challenge in 3D dynamic obstacle detection arises from the noisy depth images captured by lightweight UAV cameras, and the limited processing power of the UAV’s onboard computer makes computationally intensive methods like [35] impractical. To achieve high-accuracy results while maintaining the computational cost manageable, we propose an ensemble detection method that integrates two lightweight detectors. The first, known as the DBSCAN detector, uses the DBSCAN clustering algorithm on point cloud data from depth images to determine obstacle centers and sizes based on the boundary points in each cluster. The second, the U-depth detector, processes raw depth images to create a U-depth map (similar to a top-down view) and uses a contiguous line-grouping algorithm [16][7] to detect 3D bounding boxes of obstacles. Note that, due to the noisy input data, both detectors can produce a significant number of false-positive results. The proposed ensemble detection method addresses this issue by identifying the mutual agreements between the two detectors. In addition, since the detection results may include both static and dynamic obstacles, a lightweight YOLO detector is applied to the re-projected 2D bounding boxes in the image plane from the 3D results to classify dynamic obstacles.
@@
which includes information about the obstacle’s position, dimensions, and the length and standard deviation of the point cloud. To associate obstacles between the current and previous time steps, the two detected obstacles, $\text{o}_{i}$ and $\text{o}_{j}$, with the highest similarity score are selected, calculated as:
@@
where $W$ is a diagonal matrix that weights the obstacle features. After the association process is complete, a linear Kalman filter with a constant acceleration model is applied to estimate the positions and velocities of the obstacles.
Out-of-View Compensation: Unlike stationary sensors, the onboard camera moves with the robot during collision avoidance maneuvers. As shown in Fig. 3a, the robot may initially generate a safe trajectory with the obstacle entirely within the camera’s field of view. However, Fig. 3b illustrates that the replanned trajectory can become unsafe if the obstacle moves out of view during the maneuver. To address this issue, out-of-view compensation is introduced to estimate the states of obstacles when tracking is lost. Whenever a previously tracked dynamic obstacle moves out of view, instead of completely ignoring the obstacle, its position is linearly extrapolated using its estimated velocity, and its size is gradually inflated to account for uncertainties over a short time horizon. In our experiments, this out-of-view compensation significantly improves trajectory safety.
III-B Intent and Trajectory Prediction
The prediction module, which activates after the perception, contains two steps: the high-level intent prediction and the trajectory-level prediction. The proposed method forecasts the potential high-level actions of dynamic obstacles and then performs the trajectory prediction step to generate possible future states for all intents (Fig. 4a). This process is crucial for comprehensively anticipating the behaviors of dynamic obstacles and generating safe trajectories.
@@ III-B Intent and Trajectory Prediction
Here we define four intents to consider the forwarding, turning, and stopping decisions of dynamic obstacles. Based on the historical trajectory of the $i$-th dynamic obstacle, our goal is to estimate the probability distribution $P^{o_{i}}_{\mathcal{T}}(\mathcal{I})$ over the set of possible intents at the current time $\mathcal{T}$. Note that dynamic obstacles are assumed to move in the 2D plane.
We begin by computing the raw intent probability $\tilde{P}^{o_{i}}_{t}(\mathcal{I})$, which represents the intent probability distribution considering only the obstacle states at the time $t$ and the previous time step $t-1$. Given the obstacle states, the motion angle $\theta_{t}$ is defined as the angle between the obstacle current velocity $v_{t}$ and its facing direction vector $f_{t}=p_{t}-p_{t-1}$. For clarity, we omit the superscript for obstacle $o_{i}$ and use $p_{t}$ to represent obstacle position at time $t$. Consequently, the raw intent probability distribution $\tilde{P}^{o_{i}}_{t}$ can be computed as follows:
@@
where the values of the raw intent probability distribution is calculated based on the order of the intent set $\mathcal{I}$, with $\alpha,\beta,\gamma$ being user-defined parameters that control the probability balance across various intents. The underlying intuition is that as the motion angle goes close to zero, the probability of the forward intent increases relative to the turning intents. Similarly, when the obstacle’s speed is low, the probability of the stopping intent increases compared to the other intents.
@@
where $(\cdot\otimes\textbf{1}^{T})$ represents the operation of expanding a vector into a square matrix, and $\Sigma_{t}$ is a diagonal matrix with all entries equal to one, except for the $n$-th entry, which is set to $s(s>1)$ corresponding to the previous intent index with the highest probability. This setup of $\Sigma_{t}$ reflects the obstacle’s tendency to continue with its previous intent. The transition matrix $T^{o_{i}}_{t}$ is then obtained by normalizing each row of $\hat{T}^{o_{i}}_{t}$.
@@
Trajectory Prediction: The trajectory-level prediction computes the future positions and risk sizes of the obstacle for all possible intents, as outlined in Alg. 1. It is important to note that for each obstacle, trajectories are predicted for all four intents, regardless of the intent probability distribution. For the stop intent (Lines 1-1), the predicted positions replicate the current obstacle position, while the risk sizes are inflated based on the obstacle’s velocity and the time elapsed. For the forward, left, and right intents, a set of control inputs based on the current velocity is first generated (Lines 1-1). Constant linear and angular velocities are then applied to propagate the obstacle positions. The forward intent generates controls by varying the linear velocity within a predefined range while setting the angular velocity to zero (Line 1), whereas the left and right intents adjust both angular and linear velocities within their respective predefined ranges (Line 1). By applying these control combinations to the obstacle (Lines 1-1) with propagation continuing until a collision occurs, multiple trajectories are sampled for each intent. Finally, the predicted trajectory is determined by averaging the positions of all sampled trajectories at each time step, and the risk sizes of the obstacle are inflated by adding a value proportional to the standard deviation of the sampled trajectory positions (Lines 1-1). If the mean trajectory is not collision-free, a sampled trajectory closest to the mean trajectory will be selected. Fig. 4b visualizes the sampled trajectories for the right intent as red dotted curves and the final predicted trajectory as a blue solid curve.

### v2

@@ Abstract
Aerial robots can enhance construction site productivity by autonomously handling inspection and mapping tasks. However, ensuring safe navigation near human workers remains challenging. While navigation in static environments has been well studied, navigating dynamic environments remains open due to challenges in perception and planning. Payload limitations restrict the robots to using cameras with limited fields of view, resulting in unreliable perception and tracking during collision avoidance. Moreover, the rapidly changing conditions of dynamic environments can quickly make the generated optimal trajectory outdated.To address these challenges, this paper presents a comprehensive navigation framework that integrates perception, intent prediction, and planning. Our perception module detects and tracks dynamic obstacles efficiently and handles tracking loss and occlusion during collision avoidance. The proposed intent prediction module employs a Markov Decision Process (MDP) to forecast potential actions of dynamic obstacles with the possible future trajectories. Finally, a novel intent-based planning algorithm, leveraging model predictive control (MPC), is applied to generate navigation trajectories. Simulation and physical experiments [^3] Experiment video link: https://youtu.be/4xsEeMB9WPY demonstrate that our method improves the safety of navigation by achieving the fewest collisions compared to benchmarks.
Index Terms:
Aerial Systems: Perception and Autonomy; Integrated Planning and Control; RGB-D Perception
@@ I Introduction
Lightweight unmanned aerial vehicles (UAVs) have shown great potential for inspection and mapping in indoor construction sites [1][2][3]. Unlike open outdoor areas, indoor sites contain moving workers, machinery, and other obstacles in close proximity, making operational safety a critical concern. Consequently, a navigation framework with improved safety in these dynamic environments is essential.
Despite the success of prior research [4][5] in complex static environments, safe navigation in dynamic settings remains challenging for two main reasons. First, the low image quality of onboard cameras in lightweight UAVs makes accurate detection and tracking of dynamic obstacles difficult. Although our previous work [6] utilizes an ensemble detection method to improve accuracy and efficiency, the camera’s limited field of view (FOV) can still cause tracking failures during collision avoidance, increasing the risk of severe collisions with undetected dynamic obstacles. Second, the uncertain motion of dynamic obstacles can quickly make an optimized trajectory ineffective. While many works [7][8][9][10] incorporate obstacle trajectory prediction into planning, they generally assume a single future mode (i.e., one trajectory) per obstacle, typically using a linear motion model. While these approaches can improve dynamic obstacle avoidance, effectiveness drops significantly if the actual motion of the obstacles differs from the assumed model.
@@
To address these challenges, this paper presents an intent prediction-driven navigation framework based on a model predictive control scheme. A perception module for dynamic obstacle detection and tracking is introduced to mitigate the lost-tracking issue during collision avoidance maneuvers. Crucially, to enable the planner to consider all potential future movements of dynamic obstacles, this work utilized dynamic obstacle intents, representing high-level behaviors such as turning, stopping, or moving straight, modeled as a probability distribution. By modeling each obstacle’s intent as a probability distribution over these possible actions, we capture the uncertainties more effectively. Instead of predicting a single trajectory per obstacle, the prediction module generates all possible trajectories reflecting the various intents. The proposed planning algorithm optimizes multiple trajectories according to future states using model predictive control and applies a scoring function to select the best trajectory. Fig. 1 demonstrates our UAV navigating a dynamic environment. The contributions of this work are:
  • Intent Prediction-Driven Navigation Framework: We introduce an intent prediction-driven navigation framework comprising three key components: (1) a perception module detecting both static and dynamic obstacles with lost-tracking awareness, (2) a prediction module forecasting dynamic obstacle intents and trajectories, and (3) a planning and control module using model predictive control to generate trajectories. The framework is made available as an open-source package on GitHub  [^4] Software available at: https://github.com/Zhefan-Xu/Intent-MPC.
  • Intent and Trajectory Prediction: This work adopts a Markov Decision Process (MDP) for dynamic obstacle intent prediction to achieve high-level obstacle action forecasting and future trajectory prediction. Unlike the previous MDP-based approach [11], which relies on long-term data to model task-level intent and is therefore less suitable for real-time navigation, our method enables short-term prediction of motion-level intent.
  • Physical Flight Experiments: We validate the effectiveness of our proposed framework through extensive physical flight experiments in dynamic environments. The results demonstrate that our method enables the robot to navigate safely in various conditions.
@@ II Related Work
Reactive methods: These methods treat dynamic obstacles as static at each time step, adapting static planning methods to dynamic environments through high-frequency replanning. In [4][5], occupancy-based maps are adopted for environment representation, which enables safe trajectory generation. Ren et al. [12] achieve aggressive maneuvering using whole-body motion planning. However, these static-focused methods, lacking dynamic obstacle perception and planning, can fail in dynamic scenarios. To address the future uncertainties of dynamic obstacles, Guo et al. [13] construct risk regions, while [14] adopts a safe flight corridor that considers both static and dynamic obstacles. In [15], moving obstacle avoidance is achieved by Riemannian motion policy. Though some reactive methods enable reliable navigation, they become less effective in highly dynamic environments.
@@
Predictive methods: Unlike reactive approaches, these methods utilize obstacle future predictions for planning. In [16][7], a chance-constrained model predictive control approach is used to account for the uncertainties of dynamic obstacles. Wang et al. [8] enhance safe navigation by combining occlusion-aware obstacle tracking with spline-based trajectory optimization. Chen et al. [17] sample and evaluate trajectories based on predicted dynamic obstacles, while [10] incorporates UAV yaw angle for collision avoidance. In [9], a vision-aided planner is used for avoiding both static and dynamic obstacles. For dynamic obstacle perception, a learning-based method [18] is proposed for 3D detection and tracking. Additionally, the method proposed in [19] uses learning-based approaches to avoid dynamic obstacles. However, it relies on a constant velocity model for predicting the future state of dynamic obstacles, which may not always be reliable. To better model pedestrian trajectories, Peddi et al. [20] trained a Hidden Markov Model on datasets, and Thomas et al. [21] applied self-supervised learning for map-level dynamic obstacle prediction. Although these methods, along with other popular learning-based approaches [22][23], improve upon traditional linear predictions, they primarily focus on low-level trajectory prediction and fail to account for the high-level decision-making processes of dynamic obstacles, potentially overlooking variations in obstacle behavior. Inspired by autonomous driving methods [24][25][26][27] that predict multiple vehicle trajectories, we utilized the concept of dynamic obstacle intent in our planning framework. Our approach leverages obstacles’ high-level decisions to forecast their trajectories and generate corresponding UAV trajectories for avoidance. Unlike existing methods that aim for long-term predictions and require extensive sensing, our lightweight prediction is tailored to UAVs’ limited sensory data and is integrated into our planning algorithm to enhance the safety of navigation.
III Problem Definition
This paper addresses the problem of UAV navigation in environments with dynamic obstacles, which primarily consist of human workers whose motion can be assumed to be non-holonomic. The UAV, equipped with an RGB-D camera, must navigate such environments by following a reference trajectory while avoiding collisions with both static and dynamic obstacles in $\mathbb{R}^{3}$ space. We define safe navigation as collision-free traversal and aim to minimize collisions. The UAV system adopts a linear dynamics model (Eqn. 1) where $I$ is a 3x3 identity matrix, $\mathbf{p},\mathbf{v},\mathbf{a}\in\mathbb{R}^{3}$ represent position, velocity, and acceleration, respectively.
$\begin{bmatrix}\mathbf{p}_{k}\\ \mathbf{v}_{k}\end{bmatrix}=\begin{bmatrix}I&\Delta tI\\ 0&I\end{bmatrix}\begin{bmatrix}\mathbf{p}_{k-1}\\ \mathbf{v}_{k-1}\end{bmatrix}+\begin{bmatrix}\frac{1}{2}\Delta t^{2}I\\ \Delta tI\end{bmatrix}\mathbf{a}_{k-1}$ (1)
IV Methodology
@@
The proposed method contains three main modules: perception, prediction, and planning (Fig. 2). The perception module (Sec. IV-A) handles static and dynamic obstacles separately. The prediction module (Sec. IV-B) uses dynamic obstacle histories and a static occupancy map to generate intent probabilities and predicted trajectories. The intent-based planning algorithm (Sec. IV-D) combines a model predictive control formulation (Sec. IV-C) with a score-based trajectory selection to produce the final execution trajectory.
IV-A Lost-Tracking-Aware Perception
@@
The perception module contains static and dynamic components for obstacle perception. In Fig. 2, the static perception uses a voxel map and dynamic cleaning to remove noise from moving obstacles. The dynamic perception, based on [6], applies an updated ensemble detection method and out-of-view compensation to enhance accuracy and safety during collision avoidance, particularly in tracking loss scenarios.
@@
Figure 3: Illustration of unsafe dynamic obstacle avoidance due to tracking loss. (a) The robot initially detects the dynamic obstacle, generating a collision-free trajectory. (b) As the robot maneuvers, the obstacle moves out of the camera view, leading to an unsafe replanned trajectory.
@@
Dynamic Obstacle Detection: 3D dynamic obstacle detection is challenging due to noisy depth images from UAV cameras and limited computational resources, making methods like [28] impractical. To achieve high-accuracy results while maintaining the computational cost manageable, we propose an ensemble detection method that integrates two lightweight detectors. The first, known as the DBSCAN detector, uses the DBSCAN clustering algorithm [29] on point cloud data from depth images to determine obstacle centers and sizes based on the boundary points in each cluster. The second, the U-depth detector, processes raw depth images to create a U-depth map (similar to a top-down view) and uses a contiguous line-grouping algorithm [30][7] to detect 3D bounding boxes of obstacles. Note that, due to the noisy input data, both detectors can produce a significant number of false-positive results. The proposed ensemble detection method addresses this issue by identifying the mutual agreements between the two detectors. Because the detection results may include both static and dynamic obstacles, we first re-project the 3D bounding boxes (obtained from the mutual agreements between the DBSCAN and the U-depth detector) onto the 2D image plane. Next, we apply a lightweight YOLO detector to determine whether the re-projected 2D bounding boxes match the YOLO detections, using an IoU threshold for verification. It is worth noting that, although the single YOLO detection approach on depth camera data can generate 3D detections, the resulting outputs may still be subject to errors stemming from noisy background data and misclassifications in the 2D detection process. Therefore, adopting an ensemble detection method is necessary for a robust perception system. We refer readers to [6] for detailed detection methods we build upon.
@@
which includes information about the obstacle’s position ($\text{pos}(\text{o}_{i})$), dimensions ($\text{dim}(\text{o}_{i})$), and the length ($\text{len}(\text{o}_{i})$) and standard deviation ($\text{std}(\text{o}_{i})$) of the obstacle’s point cloud. To associate obstacles between the current and previous time steps, the two detected obstacles, $\text{o}_{i}$ and $\text{o}_{j}$, with the highest similarity score are selected and calculated as:
@@
where $W$ is a diagonal matrix that weights the obstacle features. After the association process is complete, a linear Kalman filter with a constant acceleration model is applied to estimate the positions and velocities of the obstacles. Since the YOLO detectors can only classify dynamic objects included in their training dataset, the Kalman filter’s velocity estimation serves as an alternative check for identifying dynamic objects not present in the YOLO training set.
[…]
IV-B Intent and Trajectory Prediction
@@
Here, we define four intents to consider the forwarding, turning, and stopping decisions of dynamic obstacles. Based on the historical trajectory of the $i$-th dynamic obstacle, our goal is to estimate the probability distribution $P^{o_{i}}_{\mathcal{T}}(\mathcal{I})$ over the set of possible intents at the current time $\mathcal{T}$. Note that dynamic obstacles are assumed to move in the 2D plane.
We begin by computing the raw intent probability $\tilde{P}^{o_{i}}_{t}(\mathcal{I})$, which represents the intent probability distribution considering only the obstacle states at the time $t$ and the previous time step $t-1$. Given the obstacle states, the motion angle $\theta_{t}$ is defined as the angle between the displacement vectors of consecutive time steps, representing the change in facing direction between two consecutive time steps. The raw intent probability distribution $\tilde{P}^{o_{i}}_{t}$ can be computed as:
@@
where the values of the raw intent probability distribution is calculated based on the order of the intent set $\mathcal{I}$, with $\alpha,\beta,\gamma$ being user-defined parameters that control the probability balance across various intents. The functions in Eqn. 5 are heuristically chosen to model the probability of dynamic obstacle intents. For the stop intent, we use $1-\tanh(\cdot)$ to increase the stop probability as an obstacle’s velocity approaches zero. The forward intent is modeled with an unnormalized Gaussian distribution based on motion angles which assigns higher probabilities to smaller angles that suggest straight-line movement. For the left and right intents, a sine-based function captures the increased likelihood of left or right turning with larger motion angles.
@@
where $(\cdot\otimes\textbf{1}^{T})$ denotes an outer product operation in which each element of the vector is replicated across columns to form a square matrix. Meanwhile. $\Sigma_{t}$ is a diagonal matrix with all diagonal entries set to 1, except for the $n$-th entry. Here, $n$ indicates the intent index with the highest probability at the previous time step. This $n$-th entry is assigned a user-defined weighting parameter $s$ (with $s>1$), thereby increasing the likelihood of the obstacle continuing its previously dominant intent. The transition matrix $T^{o_{i}}_{t}$ is then obtained by normalizing each row of $\hat{T}^{o_{i}}_{t}$.
@@
Trajectory Prediction: The trajectory-level prediction computes the future positions and risk sizes of the obstacle for all possible intents, as outlined in Alg. 1. It is important to note that for each obstacle, trajectories are predicted for all four intents, regardless of the intent probability distribution. For the stop intent (Lines 1-1), the predicted positions replicate the current obstacle position, while the risk sizes are inflated based on the obstacle’s velocity and the time elapsed. For the forward, left, and right intents, a set of acceleration-based control inputs is generated (Lines 1-1). Constant linear and angular acceleration are then applied to propagate the obstacle positions. The forward intent generates controls by varying the linear acceleration within a predefined range (Line 1, whereas the left and right intents adjust both angular and linear acceleration within their respective predefined ranges (Line 1).
By applying these control combinations to the obstacle (Lines 1-1) with propagation continuing until a collision occurs, multiple trajectories are sampled for each intent. Since each obstacle’s multiple intents are evaluated separately, we assume trajectories associated with a single intent remain closely clustered. Moreover, the short prediction intervals ensure these trajectories remain similar. Consequently, the predicted trajectory is determined by averaging the positions of all sampled trajectories at each time step, and the risk sizes of the obstacle are inflated by adding a value proportional to the standard deviation of the sampled trajectory positions (Lines 1-1). If the mean trajectory is not collision-free, a sampled trajectory closest to the mean trajectory will be selected. Fig. 4b visualizes the sampled trajectories for the right intent as red dotted curves and the final predicted trajectory as a blue solid curve.